# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v4

@@ Abstract
Let NN be the maximal unipotent subgroup of the simple algebraic group of type $\Phi$. It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called the coadjoint action. Such orbits play a key role in A.A. Kirillov’s orbit method. In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms. This means that we will present a set $S$ of linear forms from $\mathfrak{n}^{*}$ such that for any coadjoint orbit there exists a unique form from $S$ belonging to that orbit. The set of canonical forms will be explicitly described in terms of supports. The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
Keywords: coadjoint orbit, the orbit method, root system.
AMS subject classification: 17B08, 17B10, 17B22, 17B25, 17B30, 17B35.
1. Introduction and the main result
Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra of $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space of $\mathfrak{n}$. The group NN is also a maximal unipotent subgroup of the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup of $G$. The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g.\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$. According to the orbit method introduced by A.A. Kirillov in 1962, coadjoint orbits play the key role in the representation theory of the group NN. Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements. They are discussed in detail in the works [An95], [AN06], [Ko12], [Ko13], [Pa08], [IPa09], [Ig09], [Ig11], [Ig12], [IS21] and [IS23].
@@ 1. Introduction and the main result
$\begin{split}(\exp(x).\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
@@
$\textup{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
@@
  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N.\lambda_{1}$ and $N.\lambda_{2}$ are different.
  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N.\lambda=N.\lambda^{\prime}$.
@@
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}.\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
The columns of the matrices are numbered by all positive roots, and the rows are numbered by certain positive roots. The matrix $A_{\lambda,\gamma}$ is the matrix $B_{\lambda,\gamma}$ with one more row. Further, let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: for any $\gamma\in\textup{supp}(\lambda)$,
$\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}.$ (1)
@@
Theorem 1.1. For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
This theorem generalizes the result of the work [IS23]. This article discusses the so-called orbits associated with rook placements. By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$. It follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$. Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
@@
In [IS23] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$. Moreover, if $D$ is a non-singular rook placement, then $\lambda=f_{D,\xi}$ belongs to $S$. Indeed, if $\gamma\in D$, then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$. Hence, in the matrix $A_{\lambda,\gamma}$ the row corresponding to $\gamma$ is zero, and condition (1) is true for $\gamma\in D$. Therefore, distinct linear forms with the same support $D$ lie in different coadjoint orbits.
A result similar to Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$. In the same work [IS23], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given. By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
@@
where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves AndrГ©’s stratification for $G_{2}$:
@@
In this case $\Phi^{+}=\{\alpha,\beta,\alpha+\beta,2\alpha+\beta,3\alpha+\beta,3\alpha+2\beta\}$. Let us introduce the notation of structural constants as in the work [IS23]: $e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta}$, $e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta}$, $e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta}$, $e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}$, $e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}$.
For almost every rook placement $D$, we have $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$. The only exception is the case $D=\{\alpha+\beta,3\alpha+\beta\}$. In this case $\mathcal{O}_{D,\xi}$ is union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$. Let us write down the equations defining the orbits for $D={\alpha+\beta,3\alpha+\beta}$. A linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}$ lies in $\Omega_{D,\xi}$ if and only if it satisfies the following system:
@@
$\begin{split}\varnothing,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
The orbits of these forms correspond precisely to all the orbits whose equations are provided in [IS23] and in the present work.
Funding. The work is supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075–15–2022–287).
Acknowledgments. I thank M.V. Ignatev and A.N. Panov for helpful comments and support during the writing of this paper.
2. Description of the set of canonical forms
In this section, we will obtain an explicit description of the set $S$ to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and to demonstrate its use. At first glance, it might seem that this set does not admit a nice description. In fact, this is not the case. The set $S$ can be described very neatly in terms of supports. A list of 880 supports is presented in the Appendix.
Proposition 2.1. The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
[…]
@@ 2. Description of the set of canonical forms
[…]
@@
First, let us reformulate condition (1). The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x.\lambda)(y):=-\lambda([x,y])$. This action is called tangent.
@@
$x=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
@@
Let $\lambda\in\mathfrak{n}^{*}$, $\beta\in\Phi^{+}$. Consider the coordinates of $x.\lambda$ in the dual basis:
$(x.\lambda)(e_{\beta})=-\lambda([\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum\limits_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
@@
We can view these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$. By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$. Clearly, $\mathrm{rk}\,A_{\lambda,\gamma}=\mathrm{rk}\,B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$). In other words, $(x.\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x.\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
@@
Example 2.2. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that such a $\lambda$ lies in $S$. To visualize what the orbit of the tangent action looks like, let us refer to the table. The highlighted blue columns correspond to the roots in $D$. The linear combination $(x.\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$. It can be seen from the construction of the table that $(x.\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column. For example, the combination $(x.\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
@@
Now let us check condition (1) for all roots from $D$. As you can see from the table, $(x.\lambda)(e_{\alpha_{2}+2\alpha_{3}})=0$. Naturally, this will always be the case for the maximal root in the support. Note that $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$. Further, $(x.\lambda)(e_{\alpha_{2}})=0$, because there are no orange cells in the blue columns in the row $\alpha_{2}$; $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})=0$ for similar reasons. Finally, $(x.\lambda)(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$, which means that it is a scalar multiple of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$.
@@
It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x.\lambda)(e_{\alpha_{4}})$. This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x.\lambda)(e_{\alpha_{2}})$, $(x.\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x.\lambda)(e_{\alpha_{3}})$, $(x.\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x.\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient. The Sufficient condition below is based on this idea.
@@
Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$. Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\textup{supp}(\lambda)$. Starting from the support $D$ and the root $\gamma$, we will construct a set $\zeta_{D,\gamma}$ of positive roots. We will add roots to $\zeta_{D,\gamma}$ one by one. We add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in\zeta_{D,\gamma}$. Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots. Note that $\zeta_{D,\gamma}$ does not depend on the specific values of the coordinates of $\lambda$ in the dual basis, but only on its support $D$. The set $\zeta_{D,\gamma}$ also does not depend on the order in which we add the roots. Indeed, define sets $R_{n}$ for $n\in\mathbb{N}$. A root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in R_{i}$ for some $i<n$. Clearly, we must add all roots from $\bigcup_{n\in\mathbb{N}}R_{n}$ to $\zeta_{D,\gamma}$, and we cannot add any other roots.
@@
Proposition 2.3. (Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$. Suppose that for every root $\gamma\in D$, the following condition holds: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $\alpha\in\zeta_{D,\gamma}$. Then $\lambda\in S$.

### v5

@@ Abstract
Let NN be the maximal unipotent subgroup of the simple algebraic group of type $\Phi$. It naturally acts on the space $\mathfrak{n}^{*}$ dual to the Lie algebra $\mathfrak{n}$ of NN, and this action is called the coadjoint action. Such orbits play a key role in $A.$ Kirillov’s orbit method. In this work, we classify the orbits of this action in the case of $\Phi=F_{4}$ in terms of supports of canonical forms. This means that we will present a set $S$ of linear forms from $\mathfrak{n}^{*}$ such that for any coadjoint orbit there exists a unique form from $S$ belonging to that orbit. The set of canonical forms will be explicitly described in terms of supports. The support of a form $\lambda\in\mathfrak{n}^{*}$ is the set of positive roots $\alpha\in\Phi^{+}$ such that $\lambda(e_{\alpha})\neq 0$.
keywords:
Coadjoint orbit, the orbit method, root system.
[…]
1 Introduction and the main result
Let $\mathfrak{g}$ be a simple finite-dimensional Lie algebra over a finite field $\mathbb{F}_{q}$ of sufficiently large characteristic $p$, $\mathfrak{b}$ be a Borel subalgebra of $\mathfrak{g}$, $\Phi$ be the root system of $\mathfrak{g}$, $\Phi^{+}$ be the set of positive roots corresponding to $\mathfrak{b}$, $\mathfrak{n}$ be the nilradical of $\mathfrak{b}$, $N=\exp(\mathfrak{n})$ be the corresponding unipotent algebraic group, and $\mathfrak{n}^{*}$ be the dual space of $\mathfrak{n}$. The group NN is also a maximal unipotent subgroup of the simple algebraic group $G$ of type $\Phi$, or equivalently a Sylow $p$-subgroup of $G$. The group NN acts on $\mathfrak{n}$ by the adjoint action; the dual action of NN on the space $\mathfrak{n}^{*}$ is called the coadjoint action; we will denote the result of this action by $g\cdot\lambda$ for $g\in N$, $\lambda\in\mathfrak{n}^{*}$. According to the orbit method introduced by A.A. Kirillov in 1962, coadjoint orbits play the key role in the representation theory of the group NN. Almost all coadjoint orbits that have been studied well so far are so-called orbits associated with rook placements. They are discussed in detail in the works [1, 2, 11, 12, 13, 8, 4, 5, 7, 9, 6].
@@ 1 Introduction and the main result
$\begin{split}(\exp(x)\cdot\lambda)(y)=\lambda(\exp(-\mathrm{ad}_{x})(y))=\lambda(y)-\lambda([x,y])+\dfrac{1}{2!}\lambda([x,[x,y]])-\ldots,\\ x,y\in\mathfrak{n},~\lambda\in\mathfrak{n}^{*}.\end{split}$
@@
$\operatorname{supp}(\lambda)=\{\alpha\in\Phi^{+}\colon\lambda(e_{\alpha})\neq 0\}.$
@@
  • For any two linear forms $\lambda_{1}\neq\lambda_{2}$ whose supports belong to the given list (even if the supports coincide), the orbits $N\cdot\lambda_{1}$ and $N\cdot\lambda_{2}$ are different.
  • For any linear form $\lambda\in\mathfrak{n}^{*}$, there exists a form $\lambda^{\prime}$ with a support from the given list such that $N\cdot\lambda=N\cdot\lambda^{\prime}$.
@@
$\begin{split}A_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succeq\gamma,\beta\in\Phi^{+}}\cdot\\ B_{\lambda,\gamma}:=\left(\lambda([e_{\alpha},e_{\beta}])\right)_{\alpha\in\Phi^{+},\alpha\succ\gamma,\beta\in\Phi^{+}}.\end{split}$
The columns of the matrices are numbered by all positive roots, and the rows are numbered by certain positive roots. The matrix $A_{\lambda,\gamma}$ is the matrix $B_{\lambda,\gamma}$ with one more row. Further, let $S$ be the set of $\lambda\in\mathfrak{n}^{*}$ for which the following condition holds: for any $\gamma\in\operatorname{supp}(\lambda)$,
$\operatorname{rk}A_{\lambda,\gamma}=\operatorname{rk}B_{\lambda,\gamma}.$ (1)
@@
Theorem $1.$
For any coadjoint orbit, there exists a unique linear form $\lambda\in S$ lying on this orbit.
This theorem generalizes the result of the work [6]. This article discusses the so-called orbits associated with rook placements. By definition, a subset $D$ of $\Phi^{+}$ is called a non-singular rook placement if $\alpha-\beta\notin\Phi^{+}$ for all distinct $\alpha,\beta\in D$. It follows that $(\alpha,\beta)\leq 0$ for all distinct $\alpha,\beta\in D$. Given a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$, we put
@@
In [6] it is proved that if $\Phi=F_{4}$, $D$ is an orthogonal non-singular rook placement and $\xi_{1}$, $\xi_{2}$ are distinct maps from $D$ to $\mathbb{C}^{\times}$ then $\Omega_{D,\xi_{1}}\neq\Omega_{D,\xi_{2}}$. Moreover, if $D$ is a non-singular rook placement, then $\lambda=f_{D,\xi}$ belongs to $S$. Indeed, if $\gamma\in D$, then $\gamma+\beta\notin D$ for all $\beta\in\Phi^{+}$ and $\lambda([e_{\gamma},e_{\beta}])=0$ for all $\beta\in\Phi^{+}$. Hence, in the matrix $A_{\lambda,\gamma}$ the row corresponding to $\gamma$ is zero, and condition (1) is true for $\gamma\in D$. Therefore, distinct linear forms with the same support $D$ lie in different coadjoint orbits.
A result similar to Theorem 1.1 holds for the root system $\Phi=G_{2}$ and field $\mathbb{C}$. In the same work [6], the equations defining the basic subvarieties $\mathcal{O}_{D,\xi}$ are given. By definition, the basic subvariety $\mathcal{O}_{D,\xi}$ corresponding to a rook placement $D$ and a map $\xi\colon D\to\mathbb{C}^{\times}$ is
@@
where $\xi_{\alpha}$ is the restriction of $\xi$ to $\{\alpha\}$. That work also proves André’s stratification for $G_{2}$:
@@
In this case $\Phi^{+}=\{\alpha,\beta,\alpha+\beta,2\alpha+\beta,3\alpha+\beta,3\alpha+2\beta\}$. Let us introduce the notation of structural constants as in the work [6]:
$e_{\alpha},e_{\beta}]=c_{1}\cdot e_{\alpha+\beta},\quad[e_{\alpha},e_{\alpha+\beta}]=c_{2}\cdot e_{2\alpha+\beta},\quad[e_{\alpha},e_{2\alpha+\beta}]=c_{3}\cdot e_{3\alpha+\beta},$
$[e_{3\alpha+\beta},e_{\beta}]=c_{4}\cdot e_{3\alpha+2\beta}\quad\text{and}\quad[e_{\alpha+\beta},e_{2\alpha+\beta}]=c_{5}\cdot e_{3\alpha+2\beta}.$
For almost every rook placement $D$, we have $\mathcal{O}_{D,\xi}=\Omega_{D,\xi}$. The only exception is the case $D=\{\alpha+\beta,3\alpha+\beta\}$. In this case $\mathcal{O}_{D,\xi}$ is the union of orbits $\Omega_{D,\xi}$ and $\Omega_{\{\beta,\alpha+\beta,3\alpha+\beta\},\xi^{\prime}}$ where $\xi^{\prime}\colon{\beta,\alpha+\beta,3\alpha+\beta}\to\mathbb{C}^{\times}$ is such that its restriction to $D$ coincides with $\xi$. Let us write down the equations defining the orbits for $D={\alpha+\beta,3\alpha+\beta}$. A linear form $\lambda=\sum_{\gamma\in\Phi^{+}}\lambda_{\gamma}e_{\gamma}^{*}$ lies in $\Omega_{D,\xi}$ if and only if it satisfies the following system:
@@
$\begin{split}\emptyset,\{\alpha\},\{\beta\},\{\alpha,\beta\},\{\alpha+\beta\},\{2\alpha+\beta\},\{2\alpha+\beta,\beta\},\{3\alpha+\beta\},\\ \{3\alpha+\beta,\alpha+\beta\},\{3\alpha+\beta,\alpha+\beta,\beta\},\{3\alpha+\beta,\beta\},\{3\alpha+2\beta\},\{3\alpha+2\beta,\alpha\}.\end{split}$
The orbits of these forms correspond precisely to all the orbits whose equations are provided in [6] and in the present work.
2 Description of the set of canonical forms
In this section, we will obtain an explicit description of the set $S$ to prove Theorem 1.1 and to demonstrate its use. At first glance, it might seem that this set does not admit a nice description. In fact, this is not the case. The set $S$ can be described very neatly in terms of supports. A list of 880 supports is presented in the Appendix.
Proposition 2.1.
The set $S$ consists of all linear forms with supports from the list and forms with one of the three supports
$\begin{split}&\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2}\},\\ &\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{2},~\alpha_{4}\},\\ &\{\alpha_{1}+\alpha_{2}+2\alpha_{3},~\alpha_{1}+\alpha_{2},~\alpha_{2}+2\alpha_{3}+\alpha_{4},~\alpha_{2}+\alpha_{3}+\alpha_{4},~\alpha_{4}\},\end{split}$
@@ 2 Description of the set of canonical forms
$\alpha_{3}+\alpha_{4}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{2}$ $cellcolor[HTML]{ffa500}\alpha_{2}+\alpha_{3} \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{2}+\alpha_{3}+\alpha_{4}$
@@
$\alpha_{2}+\alpha_{3}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$ $cellcolor[HTML]{ffa500}\alpha_{3} \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{3}+\alpha_{4}$ $\alpha_{3}+2\alpha_{4}$
@@
$\alpha_{2}+2\alpha_{3}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$ $2\alpha_{4}$
$\alpha_{2}+2\alpha_{3}+\alpha_{4}$ $cellcolor[HTML]{C0C0C0}0 \cellcolor$$\cellcolor[HTML]{ffa500}\alpha_{4}$
@@
First, let us reformulate condition (1). The algebra $\mathfrak{n}$ acts on $\mathfrak{n}^{*}$ as follows: for $x,y\in\mathfrak{n}$ and $\lambda\in\mathfrak{n}^{*}$, define $(x\cdot\lambda)(y):=-\lambda([x,y])$. This action is called tangent.
@@
$x=\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha}.$
@@
Let $\lambda\in\mathfrak{n}^{*}$, $\beta\in\Phi^{+}$. Consider the coordinates of $x\cdot\lambda$ in the dual basis:
$(x\cdot\lambda)(e_{\beta})=-\lambda([\sum_{\alpha\in\Phi^{+}}x_{\alpha}e_{\alpha},e_{\beta}])=-\sum_{\alpha\in\Phi^{+}}\lambda([e_{\alpha},e_{\beta}])x_{\alpha}.$
@@
We can view these coordinates as linear combinations of variables $x_{\alpha},\alpha\in\Phi^{+}$, with constant coefficients $\lambda([e_{\alpha},e_{\beta}])$. By definition, $A_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succeq\gamma}$, and $B_{\lambda,\gamma}$ is the matrix of linear combinations $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$. Clearly, $\operatorname{rk}A_{\lambda,\gamma}=\operatorname{rk}B_{\lambda,\gamma}$ if and only if the row numbered $\gamma$ of the matrix $A_{\lambda,\gamma}$ is a linear combination of all its other rows (i.e., the rows of the matrix $B_{\lambda,\gamma}$). In other words, $(x\cdot\lambda)(e_{\gamma})$ can be expressed as a linear combination of $\{(x\cdot\lambda)(e_{\alpha})\}_{\alpha\succ\gamma}$.
Example $2.$
@@
Let $D=\operatorname{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that such a $\lambda$ lies in $S$. To visualize what the orbit of the tangent action looks like, let us refer to the table. The highlighted blue columns correspond to the roots in $D$. The linear combination $(x\cdot\lambda)(e_{\beta})$ contains only those $x_{\alpha}$ for which $\lambda([e_{\alpha},e_{\beta}])\neq 0$, which is equivalent to $\alpha+\beta\in D$. It can be seen from the construction of the table that $(x\cdot\lambda)(e_{\beta})$ contains only $x_{\alpha}$ such that the orange root $\alpha$ is in the row corresponding to $\beta$ in the highlighted column. For example, the combination $(x\cdot\lambda)(e_{\alpha_{3}})$ contains only $x_{\alpha_{4}}$ and $x_{\alpha_{2}+\alpha_{3}}$.
@@
Now let us check condition (1) for all roots from $D$. As you can see from the table, $(x\cdot\lambda)(e_{\alpha_{2}+2\alpha_{3}})=0$. Naturally, this will always be the case for the maximal root in the support. Note that $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$. Further, $(x\cdot\lambda)(e_{\alpha_{2}})=0$, because there are no orange cells in the blue columns in the row $\alpha_{2}$; $(x\cdot\lambda)(e_{\alpha_{3}+\alpha_{4}})=0$ for similar reasons. Finally, $(x\cdot\lambda)(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$, which means that it is a scalar multiple of $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$.
@@
It is clear that we did not need to know the coefficients for $x_{\alpha_{3}}$ in either $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ or $(x\cdot\lambda)(e_{\alpha_{4}})$. This is simply because $x_{\alpha_{3}}$ as a variable can be expressed in terms of $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$, and everywhere above (i.e., in $(x\cdot\lambda)(e_{\alpha_{2}})$, $(x\cdot\lambda)(e_{\alpha_{3}+\alpha_{4}})$, $(x\cdot\lambda)(e_{\alpha_{3}})$, $(x\cdot\lambda)(e_{\alpha_{4}})$) $x_{\alpha_{3}}$ can be expressed as $(x\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ multiplied by some coefficient. The Sufficient condition below is based on this idea.
@@
Consider the tangent action $\mathfrak{n}\curvearrowright\mathfrak{n}^{*}$. Let $\gamma\in\Phi^{+}$, $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$. Starting from the support $D$ and the root $\gamma$, we will construct a set $\zeta_{D,\gamma}$ of positive roots. We will add roots to $\zeta_{D,\gamma}$ one by one. We add a root $\delta$ to $\zeta_{D,\gamma}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in\zeta_{D,\gamma}$. Using this procedure, we will add to $\zeta_{D,\gamma}$ the maximum possible number of roots. Note that $\zeta_{D,\gamma}$ does not depend on the specific values of the coordinates of $\lambda$ in the dual basis, but only on its support $D$. The set $\zeta_{D,\gamma}$ also does not depend on the order in which we add the roots. Indeed, define sets $R_{n}$ for $n\in\mathbb{N}$. A root $\delta$ lies in $R_{n}$ if there exists a root $\beta\succ\gamma$ such that $\lambda([e_{\delta},e_{\beta}])\neq 0$ and if $\lambda([e_{\alpha},e_{\beta}])\neq 0$ for some other $\alpha\in\Phi^{+}$, then $\alpha\in R_{i}$ for some $i<n$. Clearly, we must add all roots from $\bigcup_{n\in\mathbb{N}}R_{n}$ to $\zeta_{D,\gamma}$, and we cannot add any other roots.
Proposition 2.3.
@@
(Sufficient condition) Let $\lambda\in\mathfrak{n}^{*}$, $D=\operatorname{supp}(\lambda)$. Suppose that for every root $\gamma\in D$, the following condition holds: if $\lambda([e_{\alpha},e_{\gamma}])\neq 0$, then $\alpha\in\zeta_{D,\gamma}$. Then $\lambda\in S$
Proof 2.4.